Additional Problem 3. If X is a continuous random variable having cdf F, then its median is defined as that value of m for which F(m) = 0.5. Find the median for random variables with the following density functions (a) f(r)-e*, x > 0 (c) f(x) 6r(1-x), 1.Additional Problem 6. Let X be a continuous random variable with pdf (a) Compute E(X), the mean of X (b) Compute Var(X), the variance of X. (c) Find an expression for Fx(r), the edf of X. (d) Calculate P(X0) (e) Compute the mean of Y, where Y (f) Find mp, the pth quantile of X X-1 X+1
Answer #1

Additional Problem 3

(a)

f(x)=e^{-x};;,;;xgeq0

herefore ;F(x)=int_{0}^{x}e^{-x}dx=1-e^{-x}

F(m)=0.5Rightarrow 1-e^{-m}=0.5Rightarrow e^{-m}=0.5Rightarrow m=-ln;0.5=0.69315

(b)

f(x) 1 , 0<x<

herefore ;F(x)=int_{0}^{x}dx=x

F(ln) 0.5 m 0.5

(c)

f(r) = 62(1-x) , 0<x<1

herefore ;F(x)=int_{0}^{x}6x(1-x)dx=6int_{0}^{x}(x-x^2)dx=3x^2-2x^3

!!!!!!!!!!!!!!!!!!!!!!!F(m)=0.5Rightarrow 3m^2-2m^3=0.5Rightarrow 4m^3-6m^2+1=0 Rightarrow (2m-1)(2m^2-2m-1)=0

Rightarrow m=0.5,,,0.5pm i  

herefore m=0.5;;;;;[ecause ;0 leq0.5pm i leq1]

.

.

Additional Problem 6

2 f(x) = (x + 1) , -1 < x < 2

(a)

!!!!!!!!!!!!!!!E(X)=int_{-1}^{2}x,f(x),dx=int_{-1}^{2}rac{2}{9}x(x+1)dx=rac{2}{9}int_{-1}^{2}(x^2+x)dx=rac{2}{9}left ( rac{14}{3}-rac{1}{6} ight )=1

(b)

2 (20 1 3 93 122

Var(X)=E(X^2)-[E(X)]^2=rac{3}{2}-1^2=rac{1}{2}

(c)

-1 9

(d)

P(X>0)=1-P(X<0)=1-F_{X}(0)=1-left ( rac{0+1}{3} ight )^2=rac{8}{9}

(e)

Y=rac{X-1}{X+1}

22 (x-1

(f)

F_{X}(m_{p})=pRightarrow left ( rac{m_{p}+1}{3} ight )^2=pRightarrow m_{p}=3sqrt{p},-1